# The Search for Squawk: Agile Modeling in Bioacoustics

### v1

@@ 1.2 Case Studies
Monitoring birds on offshore islands is critical for conservation, as many islands support unique and threatened species. Christmas Island is home to seven endemic birds, many of which are threatened or endangered, by habitat loss, invasive species, climate change and disease [southwell2024power]. Because of the remoteness of the island, these birds are difficult to survey regularly, although monitoring detectability is highly desirable both to discover declines and to quantify positive responses to conservation interventions (e.g., [tiernan2025stable]). In addition, comprehensive bird surveys of the island require at least one highly trained birdwatcher to survey for 5-6 weeks to obtain sufficient data for occupancy modeling [southwell2024power]. Inaccessibility has provided motivation for a project trialing passive acoustic monitoring, which has produced hundreds of thousands of hours of recordings. Further, difficulty accessing the island means there was almost no acoustic data available in public archives for these species. We develop classifiers for three of these low-data species using agile modeling and demonstrate the method’s ability to scale occupancy modeling to this large dataset.
@@ 2.4.1 Honeycreeper Monitoring
The classifiers were then used to estimate call density $P(\oplus)$ for each class at the 10 recording sites along the two transects following the methods described in 2.2. Adult honeycreeper singing rates are highest during the pre-breeding season, as individuals attempt to attract mates and establish or maintain territorial boundaries, then singing drops off precipitously during incubation and nestling care [Ralph_Fancy_1994]. For each class, only data collected during months of peak vocal activity were included in call density analysis (December–June for ʻAkiapōlāʻau adults, June–December for ʻAkiapōlāʻau juveniles, February–June for ʻAlawī adults, May–August for ʻAlawī juveniles). To assess the effect of elevation on call densities we ran a linear model in R (version 4.4.2; [R_2024]) for each class. Data visualization was also conducted in R.
Background on ʻAkiapōlāʻau and ʻAlawī vocalizations  Juvenile ʻAkiapōlāʻau and ʻAlawī that have fledged from their nests give frequent, distinctive begging calls that serve as contact calls or food solicitation to their parents [Pratt_2020, Lepson_2020]. ‘Akiapōlā‘au are monogamous with a long period of parental dependency (at least 4-5 months) during which time their offspring produce series of single descending chip notes [Ralph_Fancy_aki]. Adult male ‘Akiapōlā‘au sing a finch-like warbling song used in mate attraction and territory defense year round, but sing most frequently in the prebreeding season between December–February and least during August–September [Pratt_2020]. Both sexes use 2–4 note, slurred whistle calls that rise and fall in pitch year round [Pratt_2020]. The ʻAlawī have a shorter period of parental care (less than 2 months) during which juveniles produce incessant begging calls that consists of a series of stuttering, high-pitched 2–4 note clusters [VanderWerf_1998]. ‘Alawī primary song is only given by males and is a rapid, descending trill, but both males and females sometimes sing quiet and variable warbling “whisper” songs [VanderWerf_1998]. Male singing peaks from February–June corresponding with nesting season [Ralph_Fancy_1994]. ‘Alawī calls are short, ascending whistles, which can be difficult to distinguish from those of other honeycreeper species [Lepson_2020].
@@ 2.4.3 Christmas Island Bird Monitoring
This limited set of recordings was then used as a starting point to build up the labelled training set following the sampled brute-force search process described in Section 2.1. After embedding the unlabelled audio, embeddings of the labelled examples were used as search queries, using Maximum Inner Product to compare them with random subsets of the unlabelled set, generally between 5000 and 1000 hours worth, all from Deployment 1. Starting with the highest scoring, these were inspected and labelled positive or negative until at least 20 examples were found. In the case of the Christmas Island Thrush, two different call-types were identified (“chuck” and “chitter”). The two call types were queried separately, and labeled as a single class (ie, using the “Balanced Search Query” method explored in the Simulated Experiments).
@@
Detection rates were calculated for each site for each of the target species over the whole monitoring period (Figure 3). High-confidence detections were selected based on analysis of classifier scores on a sample of validated detections, and detection rate was calculated as total number of detections divided by the number of hours of audio recorded and analysed for that site.
Figure 3: WRITEME
@@ 3.1 Honeycreeper Monitoring
With the custom species-level classifiers trained for ʻAkiapōlāʻau and ʻAlawī, we ran each over unseen data. All of the 100 top scoring samples evaluated were positive detections for both the adult and juvenile ʻAlawī classes, and the ʻAkiapōlāʻau adult class (precision = 1), while 97 were positive detections of ʻAkiapōlāʻau juveniles (precision = 0.97). Classifier ROC-AUC scores derived from validation results were also high, achieving $\geq 0.8$ for all classes (3).
@@
Figure 4: Call density estimates across an elevation gradient for two Hawaiian bird species (left ʻAkiapōlāʻau, right ʻAlawī) at two life stages. Blue bars represent adult call densities, and orange bars represent juvenile call densities.
@@ 3.3 Christmas Island Bird Monitoring
Due to the large scale of audio analyzed, the embeddings were generated using high-performance computing (HPC) infrastructure, utilizing the PBS job scheduling system for distributed computation, allowing it to be completed less than one day. Once embeddings were generated, the time needed to run inference using the precomputed embeddings was very fast, around one hour on a desktop computer.
The Christmas Island Emerald Dove was detected at 77 of the 82 sites (Fig. 1a) whereas the Goshawk (Fig. 1b) and Thrush (Fig. 1c) were detected at all 82 sites throughout the monitoring period. Despite these species being widespread across the island, detection rates varied considerably among sites revealing variability in site usage. The Emerald Dove, in particular, had low detection rates and was mostly detected at sites in the north-east of the island around the town settlement.
@@ 3.4 Agile Modeling Experiments
By contrast, the Hawai‘i effort used search to obtain a fixed-size training set, and trained a single classifier. The resulting classifiers had high quality for the ʻAkiapōlāʻau ($>0.95$), but only obtained medium quality for the ʻAlawī.
@@ 4.1 Honeycreeper Monitoring
We have demonstrated that the agile modeling system can be used to quickly and efficiently assess the distribution and abundance of not only individuals of different species, but also individuals in different life stages. Our findings from Experiment B: Two Species Annotation align with abundance patterns observed in other studies that indicate ʻAkiapōlāʻau and ʻAlawī densities are increasing in high-elevation pasture areas, but are decreasing in lower-elevation closed forest areas [Kendall_2023]. Even at the most abundant sites call densities were relatively low for both species, despite restricting the dataset to only their peak pre-breeding vocalization windows. Phenology of vocalization windows for these species is poorly studied and understood, and may be shifting with climate or other ecosystem changes [Iler_2021]. This highlights another opportunity for the agile modeling system described here to fill critical knowledge gaps.
Our study design does not allow us to disentangle elevation and ecotype effects on call densities, but still serves as a proof of concept for using call density to track how specific signals may vary across covariate gradients.
It was somewhat surprising that ʻAlawī call densities not only did not track those of the adults, but also did not appear to be correlated with elevation. It is important to note that Transect B Site 4 (the highest elevation site at 2004 m) did not have recordings for August or September due to recorder failure. Lack of data for these months did not affect adult call density estimates, as neither species’ peak adult vocalization period include them. However, the juvenile call densities at this site may be skewed for both species. This particular site had a noticeable discrepancy between ʻAlawī adult and juvenile call densities, but, such a discrepancy was also observed at Transect A Site 6 (the second highest site at 1993 m). It may be that ʻAlawī adults spend considerable time within the high-elevation koa pasture areas, but prefer to nest within mid-elevation open canopy forest. Alternatively, because juveniles stay close to their nest site during the fledging period, a time when their begging calls are particularly conspicuous [VanderWerf_1998, Ralph_Fancy_aki], recorder placement near a nest may lead to higher call densities than are reflective of the study area. On the other hand, recorders within productive breeding habitat may yield deceptively low juvenile call densities if not placed close to any given nest. This effect may be greater for species with shorter periods of parental care, such as with the ʻAlawī [VanderWerf_1998], but less so for species with longer juvenile dependency, like ʻAkiapōlāʻau, as the fledglings become more mobile and vocalize farther from their nests as they develop [Pratt_2020]. With this in mind, it may be better practice to assess across a study area with multiple recorders than to use a single recording site.
@@ 4.3 Christmas Island Species ID
Despite the paucity of recordings in the training set for the embedding model, iterative active learning produced high-quality classifiers for all three target species (ROC-AUC greater than 0.95), with less than an hour of analyst time each. The Christmas Island monitoring effort plans to use these models going forward. Understanding the variation in site activity for these species of interest will help to prioritise locations where acoustic monitoring would be most effective for detecting changes in vocal activity over time.
We also found that the system worked well with the large volumes of data involved. Working with uniformly random samples of the large dataset was sufficient to obtain high classification quality. The scarcest species was the Emerald Dove, with approximately 4% call density, so any 10k hour uniformly random sample of data should contain around 400 hours of audio containing the Emerald Dove. However, for extremely scarce signals, an approximate nearest neighbor search which can work with the entire dataset may be preferable to a sampled approach.
@@ 4.4 Practical recommendations
Our simulated experiments reveal important practical insights for practitioners interested in taking advantage of the framework. Firstly, embedding function quality matters, and practitioners are better served by performing agile modeling on top of the more recent embedding functions trained on bird (and other animal) vocalization data, such as BirdNet v2.3, Perch, and SurfPerch. This is in agreement with transfer learning results presented in [ghani2023], and consequently as better embedding functions are made available their performance as measured by a (relatively simpler) transfer learning evaluation protocol should be indicative of their performance when incorporated into the agile modeling framework. Secondly, absent any prior information on abundance, a top $k$ + quantile active learning strategy is a robust and well-rounded way to decide which results to surface for human validation. Finally, incorporating expert knowledge into the search query in the form of a balanced set of query examples (in terms of vocalization types) is a simple but effective way of improving retrieval performance on minority vocalization types while maintaining strong performance on majority vocalization types.
@@ 4.5 Conclusions
In this work, we have demonstrated the broad efficacy of agile modeling for bioacoustics, building on foundational embeddings from bird classifiers. Our case studies demonstrate that this method meets the efficiency, adaptability, scalability, and quality criteria necessary to serve as a transformative ecological research and conservation tool. The extreme efficiency of model development for passive acoustic monitoring projects promises to allow the investigation of a much wider range of questions for which training data is scarce or non-existent, such as juvenile call monitoring for tracking population health and reproductive success. We have shown that detection data from novel classifiers can be easily incorporated into the ecosystem understanding, as we saw for fish sounds in coral reefs. And with Christmas Island, we see that our system can scale easily to very large datasets. Moreover, our simulated experiments reveal important practical insights for practitioners interested in taking advantage of the framework.

### v2

@@ 1.2 Case Studies
Monitoring birds on offshore islands is critical for conservation, as many islands support unique and threatened species. Christmas Island is home to seven endemic birds, many of which are threatened or endangered, by habitat loss, invasive species, climate change and disease [southwell2024power]. Because of the remoteness of the island, these birds are difficult to survey regularly, although monitoring detectability is highly desirable both to discover declines and to quantify positive responses to conservation interventions (e.g., [tiernan2025stable]). In addition, comprehensive bird surveys of the island require at least one highly trained birdwatcher to survey for 5–6 weeks to obtain sufficient data for occupancy modeling [southwell2024power]. Inaccessibility has provided motivation for a project trialing passive acoustic monitoring, which has produced hundreds of thousands of hours of recordings. Further, difficulty accessing the island means there was almost no acoustic data available in public archives for these species. We develop classifiers for three of these low-data species using agile modeling and demonstrate the method’s ability to scale occupancy modeling to this large dataset.
@@ 2.4.1 Honeycreeper Monitoring
The classifiers were then used to estimate call density $P(\oplus)$ for each class at the 10 recording sites along the two transects following the methods described in 2.2. Adult honeycreeper singing rates are highest during the pre-breeding season, as individuals attempt to attract mates and establish or maintain territorial boundaries, then singing drops off precipitously during incubation and nestling care [Ralph_Fancy_1994]. Juvenile ʻAkiapōlāʻau and ʻAlawī that have fledged from their nests give frequent, distinctive begging calls that serve as contact calls or food solicitation to their parents [Pratt_2020, Lepson_2020]. For each class, only data collected during months of peak vocal activity were included in call density analysis (December–June for ʻAkiapōlāʻau adults [Pratt_2020], June–December for ʻAkiapōlāʻau juveniles [Ralph_Fancy_aki], February–June for ʻAlawī adults [Ralph_Fancy_1994], May–August for ʻAlawī juveniles [VanderWerf_1998]). To assess the effect of elevation on call densities we ran a linear model in R (version 4.4.2; [R_2024]) for each class. Data visualization was also conducted in R.
@@ 2.4.3 Christmas Island Bird Monitoring
This limited set of recordings was then used as a starting point to build up the labelled training set following the sampled brute-force search process described in Section 2.1. After embedding the unlabelled audio, embeddings of the labelled examples were used as search queries, using Maximum Inner Product to compare them with random subsets of the unlabelled set, generally between 5000 and 1000 hours worth, all from Deployment 1. Starting with the highest scoring, these were inspected and labelled positive or negative until at least 20 examples were found. In the case of the Christmas Island Thrush, two different call-types were identified (“chuck” and “chitter”). The two call types were queried separately, and labeled as a single class (i.e., using the “Balanced Search Query” method explored in the Simulated Experiments).
@@
Detection rates were calculated for each site for each of the target species over the whole monitoring period (Figure 3). High-confidence detections were selected based on analysis of classifier scores on a sample of validated detections, and detection rate was calculated as total number of detections divided by the number of hours of audio recorded and analyzed for that site.
Figure 3: Detection rates across Christmas Island sites for the three species targeted in the study: a) Emerald Dove, b) Christmas Island Goshawk, and c) Christmas Island Thrush.
@@ 3.1 Honeycreeper Monitoring
With the custom species-level classifiers trained for ʻAkiapōlāʻau and ʻAlawī, we ran each over unseen data. All of the 100 top scoring samples evaluated were positive detections for both the adult and juvenile ʻAlawī classes, and the ʻAkiapōlāʻau adult class (precision = 1), while 97 were positive detections of ʻAkiapōlāʻau juveniles (precision = 0.97). Classifier ROC-AUC scores derived from validation results were also high, achieving $\geq 0.81$ for all classes (Table 3).
@@
Figure 4: Call density estimates across an elevation gradient for two Hawaiian bird species (left ʻAkiapōlāʻau, right ʻAlawī) at two life stages. Blue bars represent adult call densities, and orange bars represent juvenile call densities. Lines through each box represent the maximum likelihood estimate of call density.
@@ 3.3 Christmas Island Bird Monitoring
Due to the large scale of audio analyzed, the embeddings were generated using high-performance computing (HPC) infrastructure, utilizing the PBS job scheduling system for distributed computation, allowing it to be completed in less than one day. Once embeddings were generated, the time needed to run inference using the precomputed embeddings was very fast, around one hour on a desktop computer.
The Christmas Island Emerald Dove was detected at 77 of the 82 sites (Figure 3a) whereas the Goshawk (Figure 3b) and Thrush (Figure 3c) were detected at all 82 sites throughout the monitoring period. Despite these species being widespread across the island, detection rates varied considerably among sites, revealing variability in site usage. The Emerald Dove, in particular, had low detection rates and was mostly detected at sites in the north-east of the island around the town settlement.
@@ 3.4 Agile Modeling Experiments
By contrast, the Hawai‘i effort used search to obtain a fixed-size training set, and trained a single classifier. The resulting classifiers had very high quality for the ʻAkiapōlāʻau ($>0.95$), and fairly high quality for the ʻAlawī ($>0.08$).
@@ 4.1 Honeycreeper Monitoring
We have demonstrated that the agile modeling system can be used to quickly and efficiently assess the distribution and abundance of not only individuals of different species, but also individuals in different life stages. Our findings from Experiment B: Two Species Annotation align with abundance patterns observed in other studies that indicate ʻAkiapōlāʻau and ʻAlawī densities are increasing in high-elevation pasture areas, but are decreasing in lower-elevation closed forest areas [Kendall_2023]. Our study design does not allow us to disentangle elevation and ecotype effects on call densities, but still serves as a proof of concept for using call density to track how specific signals may vary across covariate gradients. Even at the most abundant sites call densities were relatively low for both species, despite restricting the dataset to only their peak pre-breeding vocalization windows. Phenology of vocalization windows for these species is poorly studied and understood, and may be shifting with climate or other ecosystem changes [Iler_2021]. This highlights another opportunity for the agile modeling system described here to fill critical knowledge gaps.
It was somewhat surprising that ʻAlawī call densities not only did not track those of the adults, but also did not appear to be correlated with elevation. It may be that ʻAlawī adults spend considerable time within the high-elevation koa pasture areas, but prefer to nest within mid-elevation open canopy forest. Alternatively, because juveniles stay close to their nest site during the fledging period, a time when their begging calls are particularly conspicuous [VanderWerf_1998, Ralph_Fancy_aki], recorder placement near a nest may lead to higher call densities than are reflective of the study area. On the other hand, recorders within productive breeding habitat may yield deceptively low juvenile call densities if not placed close to any given nest. This effect may be greater for species with shorter periods of parental care, such as with the ʻAlawī [VanderWerf_1998], but less so for species with longer juvenile dependency, like ʻAkiapōlāʻau, as the fledglings become more mobile and vocalize farther from their nests as they develop [Pratt_2020]. With this in mind, it may be better practice to assess across a study area with multiple recorders than to use a single recording site.
@@ 4.3 Christmas Island Species ID
Despite the paucity of recordings in the training set for the embedding model, iterative active learning produced high-quality classifiers for all three target species (ROC-AUC greater than 0.95), with less than an hour of analyst time each. We also found that the system worked well with the large volumes of data involved. Working with uniformly random samples of the large dataset was sufficient to obtain high classification quality. The scarcest species was the Emerald Dove, with approximately 4% call density, so any 10k hour uniformly random sample of data should contain around 400 hours of audio containing the Emerald Dove. However, for extremely scarce signals, an approximate nearest neighbor search which can work with the entire dataset may be preferable to a sampled approach.
The accuracy of the classifiers enabled us to extract data on vocal activity rates of multiple target species from a large set of daily recordings collected across Christmas Island over more than a year. This gives valuable insight into the spatial usage patterns of these threatened species which will inform monitoring decisions. The Christmas Island Emerald Dove and Christmas Island Goshawk, while widespread across the island, had very low detection rates at many sites. Selecting sites for monitoring based on highest detectability for these two species would also effectively capture other species, such as the Christmas Island Thrush, which have high detectability at most sites. In this case study we aggregated data across the entire 16-month monitoring period for a view into overall site usage. However, the high-quality output of the classifiers provides further opportunity to examine vocal activity patterns at finer temporal scales, including seasonal and diel patterns. The Christmas Island monitoring effort plans to use these models going forward to design a time- and cost-effective acoustic monitoring program that can detect declines or disappearance from sites. Understanding the variation in site activity for these species will help to prioritise locations where acoustic monitoring would be most effective for detecting changes in vocal activity over time.
@@ 4.4 Practical recommendations
Our simulated experiments reveal important practical insights for practitioners interested in taking advantage of the framework. Firstly, embedding function quality matters, and practitioners are better served by performing agile modeling on top of the more recent embedding functions trained on bird (and other animal) vocalization data, such as BirdNet v2.3, Perch, and SurfPerch. This is in agreement with transfer learning results presented in [ghani2023], and consequently, as better embedding functions are made available their performance as measured by a (relatively simpler) transfer learning evaluation protocol should be indicative of their performance when incorporated into the agile modeling framework. Secondly, absent any prior information on abundance, a top $k$ + quantile active learning strategy is a robust and well-rounded way to decide which results to surface for human validation. Finally, incorporating expert knowledge into the search query in the form of a balanced set of query examples (in terms of vocalization types) is a simple but effective way of improving retrieval performance on minority vocalization types while maintaining strong performance on majority vocalization types.
@@ 4.5 Conclusions
In this work, we have demonstrated the broad efficacy of agile modeling for bioacoustics, building on foundational embeddings from bird classifiers. Our case studies demonstrate that this method meets the efficiency, adaptability, scalability, and quality criteria necessary to serve as a transformative ecological research and conservation tool. The extreme efficiency of model development for passive acoustic monitoring projects promises to allow the investigation of a much wider range of questions for which training data is scarce or non-existent, such as juvenile call monitoring for tracking population health and reproductive success, or monitoring extremely rare birds. We have shown that detection data from novel classifiers can be easily incorporated into the ecosystem understanding, as we saw for fish sounds in coral reefs. And with Christmas Island, we see that our system can scale easily to very large datasets. Moreover, our simulated experiments reveal important practical insights for practitioners interested in taking advantage of the framework.
@@ 5 Acknowledgements
Christmas Island Case Study: Project management and collection of audio recordings was undertaken by Parks Australia Science Team with support from the Christmas Island National Park team. We thank Jess Williams, Lil Blake, Nick Macgregor, Margarita Goumas, and Brendan Tiernan at Parks Australia. We also thank Cecile Espigole and Ted Pedersen for assistance in the field. Field work was conducted under James Cook University Animal Ethics Approval number A2907. The project was supported with funding from the Australian Government under the National Environmental Science Program’s Resilient Landscapes Hub.